OpenStudy (anonymous):

how do I simply (y^9)^9

OpenStudy (anonymous):

If you take a power to a power, the exponents will multiply. Example: (x^2)^3 = x^(2*3) = x^6 Hope this helps :)

OpenStudy (anonymous):

you asked this in chemistry not math but you multiply your exponents so it should be y^81